OpenStudy (anonymous):

You've been asked to order a pizza for a party. Here's a formula for this scenario: \[x=\frac{ yz }{ a}\] x = number of people at the party y = number of slices per pizza z = number of pizzas a = number of slices each person will eat Rewrite the formula to solve for the number of pizzas you should order.

OpenStudy (hihi67):

Okay first of all do you know what variables the number of pizzas you should order are?

OpenStudy (anonymous):

Yeah, it would be the variable z.. Correct?

OpenStudy (hihi67):

Yep.

OpenStudy (hihi67):

Hey dude, sorry, I have to go now. Sorry dude.

OpenStudy (anonymous):

Hah' Ight' bro. Thanks for the help anyways.

OpenStudy (anonymous):

@AllTehMaffs @Amy143 You able to help out? :P

OpenStudy (anonymous):

Bus A and Bus B leave school on a field trip when a stopwatch reads zero. Bus A travels at a constant 50 km/h, and Bus B travels at a constant 75 km/h. How far are the two buses from school when the stopwatch reads 1.5 h?

OpenStudy (anonymous):

Lol I love you Skiller. Honest I do xD But I have no idea how to solve any of these math things you mention me in lmao

OpenStudy (anonymous):

88488, that was like speaking a foreign language to me. ._. and lol Amy ;P

OpenStudy (anonymous):

:D Just letting you know that, hon. I don't want you to beat me up if i can't help O.o lol

OpenStudy (anonymous):

@amistre64 Can you help? o_o

OpenStudy (amistre64):

z is wrapped up in multiplication and division ... unwrap it

OpenStudy (amistre64):

\[x=\frac{ yz }{ a}\] \[\frac{a}{y}x=\frac{a}{y}\frac{ y~z }{ a}\] \[\frac{a}{y}x=\frac{\cancel a}{y}\frac{ y~z }{\cancel a}\] \[\frac{a}{y}x=\frac{\cancel a}{\cancel y}\frac{ \cancel y~z }{\cancel a}\] \[\frac{a}{y}x=z\]

OpenStudy (anonymous):

Thanks so much! :)!!! Great help @amistre64